# Equation Discovery with Bayesian Spike-and-Slab Priors and Efficient Kernels

### v1

@@ Abstract
Discovering governing equations from data is important to many scientific and engineering applications. Despite promising successes, existing methods are still challenged by data sparsity as well as noise issues, both of which are ubiquitous in practice. Moreover, state-of-the-art methods lack uncertainty quantification and/or are costly in training. To overcome these limitations, we propose a novel equation discovery method based on Kernel learning and BAyesian Spike-and-Slab priors (KBASS). We use kernel regression to estimate the target function, which is flexible, expressive, and more robust to data sparsity and noises. We combine it with a Bayesian spike-and-slab prior — an ideal Bayesian sparse distribution — for effective operator selection and uncertainty quantification. We develop an expectation propagation expectation-maximization (EP-EM) algorithm for efficient posterior inference and function estimation. To overcome the computational challenge of kernel regression, we place the function values on a mesh and induce a Kronecker product construction, and we use tensor algebra methods to enable efficient computation and optimization. We show the significant advantages of KBASS on a list of benchmark ODE and PDE discovery tasks.
@@ 1 Introduction
Many scientific and engineering disciplines use differential equations to model systems of interest. These equations not only allow accurate predictions (by numerical solvers), but they also provide interpretation of the mechanism, i.e., physical laws. However, for many realistic systems, it is difficult in practice to write down a full set of partial or ordinary differential equations (PDEs/ODEs). Hence, using data-driven machine learning methods to help discover underlying physical equations is a promising (perhaps even critical, depending on the application) approach to advance our understanding of these systems and to develop powerful, reliable predictive and analysis tools.
@@
There have been several approaches for equation discovery, including: Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton et al.,, 2016) and extensions (Rudy et al.,, 2017; Schaeffer,, 2017; Zhang and Ma,, 2020; Lagergren et al.,, 2020); physics-informed neural networks (PINNs) (Raissi et al.,, 2019) with $L_{1}$ regularization (Berg and Nyström,, 2019; Both et al.,, 2021); PINN using alternating direction optimization to identify equation operators for symbolic regression (PINN-SR) (Chen et al., 2021b, ); and the most recent work (Sun et al.,, 2022) that conducts Bayesian spline learning (BSL) to estimate the target function and uses a relevant vector machine (Tipping,, 2001) to select the candidate operators.
@@
Despite the promising successes, the existing methods still have challenges with data sparsity and noise. In addition, for practical usage, they often lack uncertainty calibration and/or are costly in training. For example, SINDy uses numerical differentiation to evaluate the derivatives of candidate operators, and it can easily fail on sparsely sampled data. While PINN-based approaches avoid numerical differentiation by using neural networks (NNs) to approximate the target function, they need a careful choice of the network architectures and tedious tuning of many hyperparameters, and they have by-now well-known failure modes (Krishnapriyan et al.,, 2021). Applying differentiation operators on NNs further complicates the loss and makes the training expensive (Krishnapriyan et al.,, 2021; Mojgani et al.,, 2022). Nonetheless, PINN-based methods typically underperform in ODE discovery tasks (Sun et al.,, 2022). In addition, both the SINDy and PINN methods lack uncertainty quantification of the learned equations. Although the most recent BSL method comes with uncertainty estimation, it is quite costly to learn the spline coefficients, especially in higher dimensions.
@@
To promote the performance of data-driven equation discovery and to be more amenable for practical usage, we propose a novel equation discovery method based on Kernel learning and BAyesian Spike-and-Slab priors (KBASS). The major contributions of our work are as follows.
@@
  • Model. We introduce the KBASS method. To develop KBASS, we use kernel regression to estimate the target function and its derivatives, an approach that is flexible, expressive, and more robust to data sparsity and noises; and we then combine it with a Bayesian spike-and-slab prior to select equation operators and to estimate the operator weights. As a gold standard in Bayesian sparse learning, the spike-and-slab prior not only enjoys nice theoretical properties, such as selective shrinkage (Ishwaran and Rao,, 2005), but it also enables posterior inference of selection indicators, with which we can decide the selection results and avoid hard thresholding over the weight values. Tuning the weight threshold (as in existing methods) can be troublesome and inconvenient in practice.
@@
  • Results. We examined KBASS in discovering a list of benchmark equations, including three ODE systems, a nonlinear diffusion-reaction equation (diffusion rate $10^{-4}$), Burger’s equations with small viscosity (0.1, 0.01, 0.005), and a Kuramoto-Sivashinsky equation. We compared KBASS with state-of-the-art data-driven equation discovery methods. We observed that KBASS can use a much smaller number of data points in the presence of much more noise and still recover the equations, while the competing methods failed in most cases; and that KBASS obtained more accurate weight estimation and provided reasonable uncertainty calibration. Meanwhile, KBASS exhibits significant advantages in computational efficiency.
@@ 2 Preliminaries
Kernel Regression. Denote by ${\mathcal{H}}_{\kappa}$ the Reproducing Kernel Hilbert Space (RKHS) induced by a Mercer kernel function $\kappa(\cdot,\cdot)$, and by $\|\cdot\|_{{\mathcal{H}}_{\kappa}}$ its norm. A commonly used kernel function is the square exponential (SE) kernel, $\kappa({\bf x},{\bf x}^{\prime})=\exp(-\eta\|{\bf x}-{\bf x}^{\prime}\|^{2})$, where $\eta>0$ is the kernel parameter. Given a training dataset $\mathcal{D}=\{({\bf x}_{1},y_{1}),\ldots,({\bf x}_{N},y_{N})\}$, we use the regularized regression framework to estimate the target function, $f^{*}({\bf x})=\min_{f\in{\mathcal{H}}_{\kappa}}\sum_{n=1}^{N}L(y_{n},f({\bf x}_{n}))+\sigma^{2}\|f\|_{{\mathcal{H}}_{\kappa}},$ where the solution can be shown to take the form $f({\bf x})=\sum\nolimits_{n=1}^{N}\alpha_{n}\kappa({\bf x},{\bf x}_{n})$, and each $\alpha_{n}\in\mathbb{R}$. The learning is therefore equivalent to solving an optimization problem for the coefficients $\bm{\alpha}=[\alpha_{1},\ldots,\alpha_{N}]^{\top}$, and $\|f\|_{{\mathcal{H}}_{\kappa}}=\bm{\alpha}^{\top}{\bf K}\bm{\alpha}$, where ${\bf K}$ is an $N\times N$ kernel matrix, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Denote by ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$ the function values at the training inputs. Since ${\bf f}={\bf K}\bm{\alpha}$, we can instead model the target function as
@@
where ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and $\kappa({\bf x},{\bf X})=[\kappa({\bf x},{\bf x}_{1}),\ldots,\kappa({\bf x},{\bf x}_{N})]$. The learning amounts to minimizing the following regularized loss,
$\operatornamewithlimits{argmin}_{\bf f}\sum\nolimits_{n=1}^{N}L(y_{n},f({\bf x}_{n}))+\sigma^{2}{\bf f}^{\top}{\bf K}^{-1}{\bf f}.$ (2)
When we use the squared loss, the optimum is ${\bf f}^{*}={\bf K}({\bf K}+\sigma^{2}{\bf I})^{-1}{\bf y}$, and the function estimate is $f^{*}({\bf x})=\kappa({\bf x},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}{\bf y}$. This can also be explained in a Gaussian process (GP) regression framework and $f^{*}({\bf x})$ is the posterior mean of GP prediction (Williams and Rasmussen,, 2006).
Bayesian spike-and-slab prior. This is a powerful Bayesian sparse distribution to identify dominant patterns or signals from data (Mitchell and Beauchamp,, 1988). Take basis selection as an example, for illustration. Suppose we have $M$ bases, $\{\phi_{j}({\bf x})\}_{j=1}^{M}$, and the target function is $f({\bf x})=\sum_{j=1}^{M}w_{j}\phi_{j}({\bf x})$. To select dominant or relevant bases, the spike-and-slab prior first samples a binary selection indicator $s_{j}\in\{0,1\}$ for each basis $j$, with which to sample the basis weight $w_{j}$,
$p(s_{j})=\text{Bern}(s_{j}|\rho_{0})=\rho_{0}^{s_{j}}(1-\rho_{0})^{1-s_{j}},\quad p(w_{j}|s_{j})=s_{j}\mathcal{N}(w_{j}|0,\sigma_{0}^{2})+(1-s_{j})\delta(w_{j}),$ (3)
where $\delta(\cdot)$ is a Dirac-delta function. The selection indicator $s_{j}$ decides the prior over $w_{j}$. If $s_{j}$ is $1$, meaning that basis $j$ is selected, the weight $w_{j}$ is sampled from a flat Gaussian prior with variance $\sigma_{0}^{2}$, i.e., the slab component. If $s_{j}$ is 00, the basis $j$ is pruned, and the weight $w_{j}$ is sampled from the spike prior that concentrates on zero, i.e., the spike component. Via the selection indicators $\{s_{j}\}$, the spike-and-slab prior fulfills a selective shrinkage. For the unselected bases, the weights are directly shrunk to zero by the Dirac-delta prior $\delta(\cdot)$. For the selected ones, the weights are sampled from a flat Gaussian, which performs a mild regularization and allows the weights to be adequately estimated from data. The selective shrinkage not only is observed to be empirically better than the uniform shrinkage in $L_{1}$ regularization (Mohamed et al.,, 2012; Hernandez Lobato et al.,, 2013; Fang et al.,, 2020), but also it is critical for effective selection in terms of risk misclassification (Ishwaran and Rao,, 2005).
@@ 3 Model
To efficiently estimate the target function $u$, we construct a mesh ${\mathcal{M}}$ to cover the domain and estimate the function values at the mesh. Note that the mesh is not necessary to be evenly spaced. We can randomly sample or design the locations at each input dimension, and then construct the mesh via a Cartesian product. In doing so, we will not only be able to compute efficiently (see Section 4 for details), but also we will be able to flexibly handle different geometries in the domain, by varying the dense regions; see Fig. 8 in Appendix for an illustration. Denote by $\bm{\gamma}_{j}$ the locations at each input dimension $j$, and by $d_{j}$ the size of $\bm{\gamma}_{j}$. The mesh points are
${\mathcal{M}}=\bm{\gamma}_{1}\times\bm{\gamma}_{2}\times\bm{\gamma}_{3}=\{(t^{\prime},x_{1}^{\prime},x_{2}^{\prime})|t^{\prime}\in\bm{\gamma}_{1},x_{1}^{\prime}\in\bm{\gamma}_{2},x_{2}^{\prime}\in\bm{\gamma}_{3}\}.$
Denote the function values at ${\mathcal{M}}$ by ${\mathcal{U}}=\{u(t^{\prime},x_{1}^{\prime},x_{2}^{\prime})|(t^{\prime},x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{M}}\}$. Hence, ${\mathcal{U}}$ is a $d_{1}\times d_{2}\times d_{3}$ tensor. We use kernel regression to estimate $u(\cdot)$; and, according to (1), the target function can be modeled as a kernel interpolation from ${\mathcal{U}}$,
@@
${\partial_{{x_{1}x_{2}}}}u({\bf z})={\partial_{{x_{1}x_{2}}}}\kappa({\bf z},{\mathcal{M}})\cdot{\bf K}_{{\mathcal{M}},{\mathcal{M}}}^{-1}{\rm vec}({\mathcal{U}}),$ (6)
where ${\partial_{{x_{1}x_{2}}}}\kappa({\bf z},{\mathcal{M}})=[{\partial_{{x_{1}x_{2}}}}\kappa({\bf z},\bm{\rho}_{1}),\ldots,{\partial_{{x_{1}x_{2}}}}\kappa({\bf z},\bm{\rho}_{d})]$ and all $\{\bm{\rho}_{j}\}_{j=1}^{d}$ constitute ${\mathcal{M}}$. In this way, given the current function estimate (5), we can evaluate each operator in ${\mathcal{O}}$ on the mesh ${\mathcal{M}}$. Let us denote them by $\widehat{{\mathcal{P}}}=\{{\mathcal{P}}_{1},\ldots,{\mathcal{P}}_{A}\}$ where each is a $d_{1}\times d_{2}\times d_{3}$ tensor.
Based on $\widehat{{\mathcal{P}}}$, we propose a Bayesian sparse model to identify the operators in the PDE and to estimate the operator weights (see (4)). Specifically, for each operator $j$, we use a spike-and-slab prior to sample a selection indicator $s_{j}$ and operator weight $w_{j}$,
$p(s_{j})=\text{Bern}(s_{j}|\rho_{0})=\rho_{0}^{s_{j}}(1-\rho_{0})^{1-s_{j}},\;\;\quad p(w_{j}|s_{j})=s_{j}\mathcal{N}(w_{j}|0,\sigma_{0}^{2})+(1-s_{j})\delta(w_{j}).$ (7)
@@
where $\bm{\Phi}=[{\rm vec}({\mathcal{P}}_{1}),\ldots,{\rm vec}({\mathcal{P}}_{A})]$, and ${\bf h}$ is $\frac{\partial u}{\partial t}$ evaluated at ${\mathcal{M}}$ and has been flattened into a vector. The equation likelihood (8) measures how consistent the selected equation (by the spike-and-slab prior) is with the estimated target function at the mesh ${\mathcal{M}}$. Finally, we use a Gaussian noise model to fit the measurement data,
$p({\bf y}|{\mathcal{U}})=\mathcal{N}({\bf y}_{\text{tr}}|{\bf u}_{\text{tr}},v{\bf I})$ (9)
where ${\bf u}_{\text{tr}}$ is obtained via the kernel interpolation (5) at the training inputs and ${\bf y}_{\text{tr}}=[y_{1},\ldots,y_{N}]^{\top}$ are training outputs.
Note that a particular advantage of using the spike-and-slab prior (7) is that we can estimate the posterior of each selection indicator $p(s_{j}|\mathcal{D},{\widehat{{\bf y}}})$. We can then use it to decide if an operator should be selected, e.g., checking if $p(s_{j}=1|\mathcal{D},{\widehat{{\bf y}}})>0.5$. We never need to set a threshold over the weight values, e.g., $|w_{j}|>10^{-3}$, as popular methods (like SINDy and PINN-SR) do. Tuning a weight threshold can be troublesome and inconvenient in practice: a bigger threshold can miss important operators with small weights, e.g., in Burger’s equation with a small viscosity; a smaller one, however, can easily select false operators. Moreover, the threshold is often not general: when switching to a different problem, one has to carefully tune it from scratch.

### v2

@@ Abstract
Discovering governing equations from data is important to many scientific and engineering applications. Despite promising successes, existing methods are still challenged by data sparsity and noise issues, both of which are ubiquitous in practice. Moreover, state-of-the-art methods lack uncertainty quantification and/or are costly in training. To overcome these limitations, we propose a novel equation discovery method based on Kernel learning and BAyesian Spike-and-Slab priors (KBASS). We use kernel regression to estimate the target function, which is flexible, expressive, and more robust to data sparsity and noises. We combine it with a Bayesian spike-and-slab prior — an ideal Bayesian sparse distribution — for effective operator selection and uncertainty quantification. We develop an expectation-propagation expectation-maximization (EP-EM) algorithm for efficient posterior inference and function estimation. To overcome the computational challenge of kernel regression, we place the function values on a mesh and induce a Kronecker product construction, and we use tensor algebra to enable efficient computation and optimization. We show the advantages of KBASS on a list of benchmark ODE and PDE discovery tasks. The code is available at https://github.com/long-da/KBASS.
@@ 1 INTRODUCTION
Many scientific and engineering disciplines use differential equations to model systems of interest. These equations not only allow accurate predictions (by numerical solvers), but they also provide interpretation of the mechanism, i.e., physical laws. However, for many realistic systems, it is difficult in practice to write down a full set of partial or ordinary differential equations (PDEs/ODEs). Hence, using data-driven machine learning methods to help discover underlying physical equations is a promising approach to advance our understanding of these systems and to develop powerful, reliable predictive and analysis tools.
@@
There have been several approaches for equation discovery, including: Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton et al.,, 2016) and extensions (Rudy et al.,, 2017; Schaeffer,, 2017; Zhang and Ma,, 2020; Lagergren et al.,, 2020); physics-informed neural networks (PINNs) (Raissi et al.,, 2019) with $L_{1}$ regularization (Berg and Nyström,, 2019; Both et al.,, 2021); PINN using alternating direction optimization to identify equation operators for symbolic regression (PINN-SR) (Chen et al., 2021b, ); and the recent work (Sun et al.,, 2022) using Bayesian spline learning (BSL) to estimate the target function and relevance vector machine (Tipping,, 2001) to select the candidate operators.
@@
Despite promising successes, the existing methods still have challenges with data sparsity and noise. In addition, for practical usage, they often lack uncertainty calibration and/or are costly in training. For example, SINDy uses numerical differentiation to evaluate the derivatives of candidate operators, and it can easily fail on sparsely sampled data. While PINN-based approaches avoid numerical differentiation by using neural networks (NNs) to approximate the target function, they need a careful choice of the network architectures and tedious tuning of many hyperparameters, and they are susceptible to well-known failure modes (Krishnapriyan et al.,, 2021). Applying differentiation operators on NNs further complicates the loss and makes the training expensive (Krishnapriyan et al.,, 2021; Mojgani et al.,, 2022). Yet PINN-based methods typically underperform in ODE discovery tasks (Sun et al.,, 2022). In addition, both SINDy and PINN methods lack uncertainty quantification of the learned equations. Although the most recent BSL method comes with uncertainty estimation, it is quite costly to learn the spline coefficients, especially in higher dimensions.
@@
To promote the performance of data-driven equation discovery and to be more amenable for practical usage, we propose KBASS, a novel method based on Kernel learning and BAyesian Spike-and-Slab priors. The major contributions of our work are as follows:
@@
  • Model. We use kernel regression to estimate the target function and its derivatives, an approach that is flexible, expressive, and more robust to data sparsity and noise. We combine this with a Bayesian spike-and-slab prior to select equation operators and to estimate the operator weights. As a gold standard in Bayesian sparse learning, the spike-and-slab prior has nice theoretical properties, such as selective shrinkage (Ishwaran and Rao,, 2005). It also enables posterior inference of selection indicators, with which we can decide the selection results and avoid hard thresholding over the weight values. Tuning the weight threshold (as in existing methods) can be troublesome and inconvenient in practice.
@@
  • Results. We examine KBASS in discovering a number of benchmark equations, including three ODE systems, a nonlinear diffusion-reaction equation (diffusion rate $10^{-4}$), Burgers’ equations with small viscosity (0.1, 0.01, 0.005), and a Kuramoto-Sivashinsky equation. We compared KBASS with state-of-the-art data-driven equation discovery methods. We observed that KBASS can use a much smaller number of data points in the presence of much more noise and still recover the equations, while the competing methods failed in most cases; and that KBASS obtained more accurate weight estimation and provided reasonable uncertainty calibration. Meanwhile, KBASS exhibits significant advantages in computational efficiency.
@@ 2 PRELIMINARIES
Kernel Regression. Denote by ${\mathcal{H}}_{\kappa}$ the Reproducing Kernel Hilbert Space (RKHS) induced by a Mercer kernel function $\kappa(\cdot,\cdot)$, and by $\|\cdot\|_{{\mathcal{H}}_{\kappa}}$ its norm. Given a training dataset $\mathcal{D}=\{({\bf x}_{1},y_{1}),\ldots,({\bf x}_{N},y_{N})\}$, we use the regularized regression framework to estimate the target function, $f^{*}({\bf x})=\min_{f\in{\mathcal{H}}_{\kappa}}\sum_{n=1}^{N}L(y_{n},f({\bf x}_{n}))+\sigma^{2}\|f\|_{{\mathcal{H}}_{\kappa}},$ where the solution can be shown to take the form $f({\bf x})=\sum\nolimits_{n=1}^{N}\alpha_{n}\kappa({\bf x},{\bf x}_{n})$ and each $\alpha_{n}\in\mathbb{R}$. The learning is therefore equivalent to solving an optimization problem for the coefficients $\bm{\alpha}=[\alpha_{1},\ldots,\alpha_{N}]^{\top}$, and $\|f\|_{{\mathcal{H}}_{\kappa}}=\bm{\alpha}^{\top}{\bf K}\bm{\alpha}$, where ${\bf K}$ is an $N\times N$ kernel matrix, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Denote by ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$ the function values at the training inputs. Since ${\bf f}={\bf K}\bm{\alpha}$, we can instead model the target function as
@@
where ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and $\kappa({\bf x},{\bf X})=[\kappa({\bf x},{\bf x}_{1}),\ldots,\kappa({\bf x},{\bf x}_{N})]$. The learning amounts to minimizing the following regularized loss, $\operatornamewithlimits{argmin}_{\bf f}\sum\nolimits_{n=1}^{N}L(y_{n},f({\bf x}_{n}))+\sigma^{2}{\bf f}^{\top}{\bf K}^{-1}{\bf f}$. When we use the squared loss, the optimum is ${\bf f}^{*}={\bf K}({\bf K}+\sigma^{2}{\bf I})^{-1}{\bf y}$, and the function estimate is $f^{*}({\bf x})=\kappa({\bf x},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}{\bf y}$. This can also be explained in a Gaussian process (GP) regression framework, where $f^{*}({\bf x})$ is the posterior mean of GP prediction (Williams and Rasmussen,, 2006).
Bayesian spike-and-slab prior is a powerful Bayesian sparse distribution to identify dominant patterns or signals from data (Mitchell and Beauchamp,, 1988). Take basis selection as an example for illustration. Suppose we have $M$ basis functions, $\{\phi_{j}({\bf x})\}_{j=1}^{M}$, and the target function is $f({\bf x})=\sum_{j=1}^{M}w_{j}\phi_{j}({\bf x})$. To select dominant or relevant bases, the spike-and-slab prior first samples a binary selection indicator $s_{j}\in\{0,1\}$ for each basis $j$, with which to sample the basis weight $w_{j}$,
$p(s_{j})$ $=\text{Bern}(s_{j}|\rho_{0})=\rho_{0}^{s_{j}}(1-\rho_{0})^{1-s_{j}},$
$p(w_{j}|s_{j})$ $=s_{j}\mathcal{N}(w_{j}|0,\sigma_{0}^{2})+(1-s_{j})\delta(w_{j}),$ (2)
where $\delta(\cdot)$ is a Dirac-delta function. The selection indicator $s_{j}$ decides the prior over $w_{j}$. If $s_{j}$ is $1$, meaning that basis $j$ is selected, the weight $w_{j}$ is sampled from a flat Gaussian prior with variance $\sigma_{0}^{2}$, i.e., the slab component. If $s_{j}$ is 00, the basis $j$ is pruned, and the weight $w_{j}$ is sampled from the spike prior that concentrates on zero, i.e., the spike component. Via the selection indicators $\{s_{j}\}$, the spike-and-slab prior fulfills a selective shrinkage. For the unselected bases, the weights are directly shrunk to zero by the Dirac-delta prior $\delta(\cdot)$. For the selected ones, the weights are sampled from a flat Gaussian, which performs a mild regularization and allows the weights to be adequately estimated from data. Not only is the selective shrinkage observed empirically to be better than the uniform shrinkage in $L_{1}$ regularization (Mohamed et al.,, 2012; Fang et al.,, 2020), but it is critical for effective selection in terms of risk misclassification (Ishwaran and Rao,, 2005).
@@ 3 MODEL
To efficiently estimate $u$, we construct a mesh ${\mathcal{M}}$ to cover the domain and estimate the function values at the mesh. Note that the mesh is not necessary to be evenly spaced. We can randomly sample or design the locations at each input dimension, and then construct the mesh via a Cartesian product. In doing so, we will not only be able to compute efficiently (see Section 4 for details), but also we will be able to flexibly handle different geometries in the domain by varying the dense regions; see Appendix Fig. 8 for an illustration. Denote by $\bm{\gamma}_{j}$ the locations at each input dimension $j$, and by $d_{j}$ the size of $\bm{\gamma}_{j}$. The mesh points are ${\mathcal{M}}=\bm{\gamma}_{1}\times\bm{\gamma}_{2}\times\bm{\gamma}_{3}=\{(t^{\prime},x_{1}^{\prime},x_{2}^{\prime})|t^{\prime}\in\bm{\gamma}_{1},x_{1}^{\prime}\in\bm{\gamma}_{2},x_{2}^{\prime}\in\bm{\gamma}_{3}\}$.
Denote the function values at ${\mathcal{M}}$ by ${\mathcal{U}}=\{u(t^{\prime},x_{1}^{\prime},x_{2}^{\prime})|(t^{\prime},x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{M}}\}$. Hence, ${\mathcal{U}}$ is a $d_{1}\times d_{2}\times d_{3}$ tensor. We use kernel regression to estimate $u(\cdot)$, and, according to (1), the target function can be modeled as a kernel interpolation from ${\mathcal{U}}$,
@@
${\partial_{{x_{1}x_{2}}}}u({\bf z})={\partial_{{x_{1}x_{2}}}}\kappa({\bf z},{\mathcal{M}})\cdot{\bf K}_{{\mathcal{M}},{\mathcal{M}}}^{-1}{\rm vec}({\mathcal{U}}).$ (5)
Here ${\partial_{{x_{1}x_{2}}}}\kappa({\bf z},{\mathcal{M}})\overset{\Delta}{=}[{\partial_{{x_{1}x_{2}}}}\kappa({\bf z},\bm{\rho}_{1}),\ldots,{\partial_{{x_{1}x_{2}}}}\kappa({\bf z},\bm{\rho}_{d})]$, and all $\{\bm{\rho}_{j}\}_{j=1}^{d}$ constitute ${\mathcal{M}}$. In this way, given the current function estimate (4), we can evaluate each operator in ${\mathcal{O}}$ on mesh ${\mathcal{M}}$. We denote them by $\widehat{{\mathcal{P}}}=\{{\mathcal{P}}_{1},\ldots,{\mathcal{P}}_{A}\}$ where each is a $d_{1}\times d_{2}\times d_{3}$ tensor.
Based on $\widehat{{\mathcal{P}}}$, we propose a Bayesian sparse model to identify the operators in the PDE and to estimate the operator weights. Specifically, for each operator $j$, we use a spike-and-slab prior to sample a selection indicator $s_{j}$ and operator weight $w_{j}$ as in (2).
@@
where $\bm{\Phi}=[{\rm vec}({\mathcal{P}}_{1}),\ldots,{\rm vec}({\mathcal{P}}_{A})]$, ${\bf h}$ is $\frac{\partial u}{\partial t}$ evaluated at ${\mathcal{M}}$ and has been flattened into a vector. The equation likelihood (6) measures how consistent the selected equation (by the spike-and-slab prior) is with the estimated target function at the mesh ${\mathcal{M}}$. Finally, we use a Gaussian noise model to fit the measurement data,
$p({\bf y}_{\text{tr}}|{\mathcal{U}})=\mathcal{N}({\bf y}_{\text{tr}}|{\bf u}_{\text{tr}},v{\bf I}),$ (7)
where ${\bf u}_{\text{tr}}$ is obtained via the kernel interpolation (4) at the training inputs and ${\bf y}_{\text{tr}}=[y_{1},\ldots,y_{N}]^{\top}$ are the training outputs.
Note that a particular advantage of using the spike-and-slab prior (2) is that we can estimate the posterior of each selection indicator $p(s_{j}|\mathcal{D},{\widehat{{\bf y}}})$, and use it to decide if an operator should be selected, e.g., checking if $p(s_{j}=1|\mathcal{D},{\widehat{{\bf y}}})>0.5$. We never need to set a threshold over the weight values, e.g., $|w_{j}|>10^{-3}$, as popular methods (like SINDy and PINN-SR) do. Tuning a weight threshold can be troublesome and inconvenient in practice. A bigger threshold can miss important operators with small weights, e.g., in Burger’s equation with a small viscosity; a smaller one, however, can easily select false operators. Moreover, the threshold is often not general. When switching to a different problem, one has to carefully tune it from scratch.